# Variational Continual Test-Time Adaptation

### v1

@@ 4.1 Mean-Teacher architecture
where $d$ denotes each dimension of the parameter. The teacher model $\bar{p}(\boldsymbol{\theta})$ is also a Gaussian distribution. Thus, the student model is updated by aligning it with the teacher model through the use of a cross-entropy (CE) loss
$L_{\text{CE}}(q,\bar{p})=-\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta})}\bar{p}(\mathcal{U}|\boldsymbol{\theta})\log q(\mathcal{U}|\boldsymbol{\theta}).$ (8)
In our implementation, we also try to use Symmetric Cross-Entropy (SCE) (Wang et al., 2019) in CTTA,
$L_{\text{SCE}}(q,\bar{p})=-\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta})}\big{[}\bar{p}(\mathcal{U}|\boldsymbol{\theta})\log q(\mathcal{U}|\boldsymbol{\theta})$ (9)
$+q(\mathcal{U}|\boldsymbol{\theta})\log\bar{p}(\mathcal{U}|\boldsymbol{\theta})\big{]}.$
SCE keeps the gradient for high and low confidence balanced, benefiting the optimization in unsupervised learning.
4.2 Mixture-of-Gaussian Prior
For the KL term, to reduce prior drift in unsupervised testing time, we propose a mixing-up approach to combining the teacher and source prior adaptatively. The source prior is warmed up upon the pretrained deterministic model $p_{1}(\boldsymbol{\theta})=p(\boldsymbol{\theta}|\mathcal{D}_{0})$ (see Sec. 4.3.1). The teacher model $\bar{p}_{t}(\boldsymbol{\theta})$ is updated by EMA (see Sec. 4.3.3). We assume that the prior should be the mixture of the two Gaussian priors. Using only the source prior, the adaptation is limited. While using only the teacher prior, the prior is prone to drift.
We use the mean entropy derived from a given serious data augmentation to represent the confidence of the two prior models, and mix up the two priors with a modulating factor
$\alpha=\frac{1}{|\mathcal{I}|}\sum\nolimits_{i\in\mathcal{I}}\frac{e^{H(\mathcal{U}_{i}|\boldsymbol{\theta}_{0})/\tau}}{e^{H(\mathcal{U}_{i}|\boldsymbol{\theta}_{0})/\tau}+e^{H(\mathcal{U}_{i}|\bar{\boldsymbol{\theta}})/\tau}},$ (10)
where $\mathcal{I}$ denotes augmentation types. $\boldsymbol{\theta}_{0}$ and $\bar{\boldsymbol{\theta}}$ are the parameters of the source model and the teacher model. $\tau$ means the temperature factor. Thus, as shown in Fig. 2(b), the current prior $p_{t}(\boldsymbol{\theta})$ is set to the mixture of priors as
@@ 4.2 Mixture-of-Gaussian Prior
In the VI, we use the upper bound to update the KL term (Liu et al., 2019) (see Appendix C) for simplicity,
@@
where, for brevity, we let $\bar{p}(x^{\prime}_{i})=\bar{p}(x^{\prime}_{i}|\boldsymbol{\theta})$ and $\bar{p}(x)=\bar{p}(x)|\boldsymbol{\theta})$ in short. $f(\cdot)$ is the confidence function. $\epsilon$ denotes the confidence margin and $\mathbf{1}(\cdot)$ is an indicator function. Eq. (13) means that the reliable teacher is represented by the average of its augmentations with $\epsilon$ more confidence.
@@
  Input: Source data $\mathcal{D}_{0}$, pretrained model $p_{0}(\boldsymbol{\theta})$
  Initialize prior distribution $p(\boldsymbol{\theta})$ with $p_{0}(\boldsymbol{\theta})$
  Update $p(\boldsymbol{\theta}|\mathcal{D}_{0})\approx q_{0}(\boldsymbol{\theta})$ by $p(\boldsymbol{\theta})$ and $\mathcal{D}_{0}$ using Eq. (14)
  Output: Source prior $p_{1}(\boldsymbol{\theta})=p(\boldsymbol{\theta}|\mathcal{D}_{0})$
4.3 Adaptation and Inference
4.3.1 Variational Warm-up
To obtain a source BNN, we choose not to train a model on the source data $\mathcal{D}_{0}$ from scratch. Warm-up has been used in CTTA to further build knowledge structure for the source model (Döbler et al., 2023; Brahma & Rai, 2023). Inspired by this, we design to transform the pretrained deterministic model to a BNN by variational warm-up. Specifically, we leverage the local reparameterization trick (Kingma et al., 2015) to add stochastic parameters, and warm up the model:
where $p(\boldsymbol{\theta})$ represents the prior distribution, say the pretrained deterministic model. By the variational warm-up, we can easily transform an off-the-shelf pretrained model into a BNN with a stochastic dynamic. The variational warm-up strategy is outlined in Algorithm 1.
Algorithm 2 Variational CTTA
  Input: Source data $\mathcal{D}_{0}$, pretrained model $p_{0}(\boldsymbol{\theta})$, Unlabeled test data from different domain $\mathcal{U}_{1:T}$
  $p_{1}(\boldsymbol{\theta})=\text{Variational warm-up}(\mathcal{D}_{0},p_{0}(\boldsymbol{\theta}))$. // Alg. 1
  for Domain shift $t=1$ to $T$ do
     for Test data $x\sim\mathcal{U}_{t}$ do
        Model predict for $x$ (Eq. (17))
        Update student model using $x$ (Eq. (15))
        Update teacher model via EMA (Eq. (16))
     end for
  end for
Table 1: Classification error rate (%) for the standard CIFAR10-to-CIFAR10C continual test-time adaptation task. All results are evaluated with the largest corruption severity level 5 in an online fashion.
Method C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 C11 C12 C13 C14 C15 Avg
Source 72.3 65.7 72.9 46.9 54.3 34.8 42.0 25.1 41.3 26.0 9.3 46.7 26.6 58.5 30.3 43.5
BN 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
Tent (Wang et al., 2020) 24.8 20.6 28.5 15.1 31.7 17.0 15.6 18.3 18.3 18.1 11.0 16.8 23.9 18.6 23.9 20.1
CoTTA (Wang et al., 2022) 24.5 21.5 25.9 12.0 27.7 12.2 10.7 15.0 14.1 12.7 7.6 11.0 18.5 13.6 17.7 16.3
PETAL (Brahma & Rai, 2023) 23.7 21.4 26.3 11.8 28.8 12.4 10.4 14.8 13.9 12.6 7.4 10.6 18.3 13.1 17.1 16.2
SATA (Chakrabarty et al., 2023) 23.9 20.1 28.0 11.6 27.4 12.6 10.2 14.1 13.2 12.2 7.4 10.3 19.1 13.3 18.5 16.1
SWA (Yang et al., 2023) 23.9 20.5 24.5 11.2 26.3 11.8 10.1 14.0 12.7 11.5 7.6 9.5 17.6 12.0 15.8 15.3
VCoTTA (Ours) 18.1 14.9 22.0 9.7 22.6 11.0 9.5 11.4 10.6 10.5 6.5 9.4 15.6 11.0 14.5 13.1
4.3.2 Student update
The student model $q_{t}(\boldsymbol{\theta})$ is adapted and approximated by VI in Eq. (5), the total process can be optimized by minimizing
$L(q_{t})$ $=L_{\text{SCE}}(q_{t},\bar{p}^{\prime}_{t})$ (15)
$+\alpha\cdot\text{KL}\left(q_{t}||q_{0}\right)+(1-\alpha)\cdot\text{KL}\left(q_{t}||\bar{q}_{t}\right),$
@@ 4.3.2 Student update
4.3.3 Teacher update via EMA
The teacher model is updated by EMA. Let $(\boldsymbol{\mu},\boldsymbol{\sigma})$ and $(\bar{\boldsymbol{\mu}},\bar{\boldsymbol{\sigma}})$ be the mean and standard deviation (std) of the student and teacher model, respectively. At test time, the teacher model $\bar{q}_{t}(\boldsymbol{\theta})$ is updated in real-time by
@@ 4.3.4 Model inference
[…]
@@
4.3.5 The algorithm
@@
We illustrate the whole algorithm in Algorithm 2. We first transform an off-the-shelf pretrained model into BNN via the variational warm-up strategy (Sec. 4.3.1). After that, we obtain a BNN, and for each domain shift, we forward and adapt each test data point in an MT architecture. For a data point $x$, we first predict the class label using the mixture of the source model and the teacher model (Sec. 4.3.4). Then, we update the student model using VI, where we use cross entropy to compute the entropy term and use the mixture of priors for the KL term (Sec. 4.3.2). Finally, we update the BNN teacher model via EMA (Sec. 4.3.3). See more details in Appendix D. The process is feasible for any test data without labels.
Table 2: Classification error rate (%) for the standard CIFAR100-to-CIFAR100C continual test-time adaptation task. All results are evaluated with the largest corruption severity level 5 in an online fashion. C1 to C15 are 15 corruptions for the datasets (see Sec. 5.1).
Method C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 C11 C12 C13 C14 C15 Avg
Source 73.0 68.0 39.4 29.3 54.1 30.8 28.8 39.5 45.8 50.3 29.5 55.1 37.2 74.7 41.2 46.4
BN 42.1 40.7 42.7 27.6 41.9 29.7 27.9 34.9 35 41.5 26.5 30.3 35.7 32.9 41.2 35.4
Tent (Wang et al., 2020) 37.2 35.8 41.7 37.9 51.2 48.3 48.5 58.4 63.7 71.1 70.4 82.3 88.0 88.5 90.4 60.9
CoTTA (Wang et al., 2022) 40.1 37.7 39.7 26.9 38.0 27.9 26.4 32.8 31.8 40.3 24.7 26.9 32.5 28.3 33.5 32.5
PETAL (Brahma & Rai, 2023) 38.3 36.4 38.6 25.9 36.8 27.3 25.4 32.0 30.8 38.7 24.4 26.4 31.5 26.9 32.5 31.5
SATA (Chakrabarty et al., 2023) 36.5 33.1 35.1 25.9 34.9 27.7 25.4 29.5 29.9 33.1 23.6 26.7 31.9 27.5 35.2 30.3
SWA (Yang et al., 2023) 39.4 36.4 37.4 25.0 36.0 26.6 25.0 29.1 28.4 35.0 23.5 25.1 28.5 25.8 29.6 30.0
VCoTTA (Ours) 35.3 32.8 38.9 23.8 34.6 25.5 23.2 27.5 26.7 30.4 22.1 23.0 28.1 24.2 30.4 28.4
Table 3: Classification error rate (%) for the standard ImageNet-to-ImageNetC continual test-time adaptation task. All results are evaluated with the largest corruption severity level 5 in an online fashion. C1 to C15 are 15 corruptions for the datasets (see Sec. 5.1).
Method C1 C2 C3 C4 C5 C6 C7 C8 C9 C10 C11 C12 C13 C14 C15 Avg
Source 95.3 95.0 95.3 86.1 91.9 87.4 77.9 85.1 79.9 79.0 45.4 96.2 86.6 77.5 66.1 83.0
BN 87.7 87.4 87.8 88.0 87.7 78.3 63.9 67.4 70.3 54.7 36.4 88.7 58.0 56.6 67.0 72.0
Tent (Wang et al., 2020) 85.6 79.9 78.3 82.0 79.5 71.4 59.5 65.8 66.4 55.2 40.4 80.4 55.6 53.5 59.3 67.5
CoTTA (Wang et al., 2022) 87.4 86.0 84.5 85.9 83.9 74.3 62.6 63.2 63.6 51.9 38.4 72.7 50.4 45.4 50.2 66.7
PETAL (Brahma & Rai, 2023) 87.4 85.8 84.4 85.0 83.9 74.4 63.1 63.5 64.0 52.4 40.0 74.0 51.7 45.2 51.0 67.1
VCoTTA (Ours) 81.8 78.9 80.0 83.4 81.4 70.8 60.3 61.1 61.7 46.4 35.7 71.7 50.1 47.1 52.9 64.2
5 Experiment
5.1 Experimental setting

### v2

@@ IV-A Entropy term: Integrating Mean Teacher into VI
where $d$ denotes each dimension of the parameter. The teacher model $\bar{p}(\boldsymbol{\theta})$ is also modeled as a Gaussian distribution and serves as one component of the prior mixture. Importantly, $\bar{p}(\boldsymbol{\theta})$ itself is not guaranteed to be reliable. Instead, our Bayesian formulation adaptively weights the contribution of the teacher and the source prior according to uncertainty. In practice, this adaptive weighting is realized by guiding the student distribution to align with the teacher through a cross-entropy (CE) loss:
$L_{\text{CE}}(q,\bar{p})=-\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta})}\mathbb{E}_{x\sim\mathcal{U}}\left[\bar{p}(x|\boldsymbol{\theta})\log q(x|\boldsymbol{\theta})\right].$ (9)
In practice, following [wang2019symmetric], we also experiment with the Symmetric Cross-Entropy (SCE) loss to improve training stability:
$L_{\text{SCE}}(q,\bar{p})=-\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta})}\mathbb{E}_{x\sim\mathcal{U}}\big[\bar{p}(x|\boldsymbol{\theta})\log q(x|\boldsymbol{\theta})$ (10)
$+q(x|\boldsymbol{\theta})\log\bar{p}(x|\boldsymbol{\theta})\big].$
SCE is not our contribution, but an auxiliary choice that balances gradients for high- and low-confidence samples.
IV-B KL term: Mixture-of-Gaussian Prior
For the KL term, to reduce the impact of unreliable prior, we propose a mixing-up approach to combining the teacher and source prior adaptatively. The source prior is warmed up upon the pretrained deterministic model $p_{1}(\boldsymbol{\theta})=p(\boldsymbol{\theta}|\mathcal{D}_{0})$ (see Section IV-C1). The teacher model $\bar{p}_{t}(\boldsymbol{\theta})$ is updated by EMA (see Section IV-C3). We assume that the prior should be the mixture of the two Gaussian priors. Using only the source prior, the adaptation is limited. While using only the teacher prior, the prior is prone to be unreliable.
We use the mean entropy derived from a given series of data augmentation to represent the confidence of the two prior models and mix up the two priors with a modulating factor
$\alpha=\frac{1}{|\mathcal{I}|}\sum\nolimits_{i\in\mathcal{I}}\frac{e^{H(x|\boldsymbol{\theta}_{0})/\tau}}{e^{H(x|\boldsymbol{\theta}_{0})/\tau}+e^{H(x|\bar{\boldsymbol{\theta}})/\tau}},$ (11)
where $\mathcal{I}$ denotes augmentation types. $\boldsymbol{\theta}_{0}$ and $\bar{\boldsymbol{\theta}}$ are the parameters of the source model and the teacher model. $\tau$ is the temperature factor. Thus, as shown in Fig. 2(b), the current prior $p_{t}(\boldsymbol{\theta})$ is set to the mixture of priors as
@@ IV-B KL term: Mixture-of-Gaussian Prior
In the VI, we use the upper bound to update the KL term [liu2019variational] (see Section V-B) for simplicity,
@@
where, for brevity, we let $\bar{p}(x^{\prime}_{i})=\bar{p}(x^{\prime}_{i}|\boldsymbol{\theta})$ and $\bar{p}(x)=\bar{p}(x)|\boldsymbol{\theta})$ in short. $f(\cdot)$ is the confidence function. $\epsilon$ denotes the confidence margin and $\mathbf{1}(\cdot)$ is an indicator function. Eq. (14) can be regarded as a filter, meaning that for each sample, the reliable teacher is represented by the average of its augmentations with $\epsilon$ more confidence. In Section V-C, we prove that the proposed mixture-of-Gaussian is benifical to CTTA. In the experiments of Section VI-G1, we discuss the influence of different $\epsilon$.
IV-C Adaptation and Inference
IV-C1 Variational Warm-up
To obtain a source BNN, instead of training a model from scratch on the source data $\mathcal{D}_{0}$, we transform a pretrained deterministic model to a BNN by variational warm-up strategy. Specifically, we leverage the local reparameterization trick [kingma2015variational] to add stochastic parameters, and warm up the model:
where $p(\boldsymbol{\theta})$ represents the prior distribution, say the pretrained deterministic model. Eq. (15) denotes a standard VI on the source data, and we optimize the ELBO to obtain the variational distribution [wang2011online]. By the variational warm-up, we can easily transform an off-the-shelf pretrained model into a BNN with a stochastic dynamic. The variational warm-up strategy is outlined in Algorithm 1.
@@ IV-C1 Variational Warm-up
1:Input: Source data $\mathcal{D}_{0}$, pretrained model $p_{0}(\boldsymbol{\theta})$
2: Initialize prior distribution $p(\boldsymbol{\theta})$ with $p_{0}(\boldsymbol{\theta})$
3: Update $p(\boldsymbol{\theta}|\mathcal{D}_{0})\approx q_{0}(\boldsymbol{\theta})$ by $p(\boldsymbol{\theta})$ and $\mathcal{D}_{0}$ using Eq. (15)
4:Output: Source prior $p_{1}(\boldsymbol{\theta})=p(\boldsymbol{\theta}|\mathcal{D}_{0})$
The warm-up strategy is a common approach in TTA and CTTA tasks to further build knowledge structure for the source model, such as [jung2023cafa, song2023ecotta, dobler2023robust, chen2023each]. Some other methods may not use warm-up but still use the source data, such as [niu2022efficient]. The warm-up strategy uses the source data only before deploying the model to CTTA scenario, and it is regarded as a part of pretraining. All these methods using source data are operationalized source-free at test time and found beneficial to CTTA. We use the warm-up to inject the uncertainties into a given source model, i.e., turning an off-the-shelf pretrained model into a pretrained BNN model. This is convenient to obtain a pretrained BNN, because the warm-up strategy uses only a few epochs. We offer more discussions and experiments on the proposed variational warm-up strategy in Section VI-H.
IV-C2 Student update via VI
The student model $q_{t}(\boldsymbol{\theta})$ is adapted by approximating using Eq. (5), and is optimized on:
$L(q_{t})=L_{\text{CE}}(q_{t},\bar{p}^{\prime}_{t})+\alpha\cdot\text{KL}\left(q_{t}||q_{0}\right)+(1-\alpha)\cdot\text{KL}\left(q_{t}||\bar{q}_{t}\right),$ (16)
@@ IV-C2 Student update via VI
IV-C3 Teacher update via EMA
The teacher model is updated using EMA. Let $(\boldsymbol{\mu},\boldsymbol{\sigma})$ and $(\bar{\boldsymbol{\mu}},\bar{\boldsymbol{\sigma}})$ be the mean and standard deviation of the student and teacher model, respectively. At test time, the teacher model $\bar{q}_{t}(\boldsymbol{\theta})$ is updated by
@@ IV-C4 Model inference
$p_{t}(x)=$ $\int p(x|\boldsymbol{\theta})p_{t}(\boldsymbol{\theta})d\boldsymbol{\theta}$ (18)
@@
IV-D The algorithm
Algorithm 2 Variational CTTA
1:Input: Source data $\mathcal{D}_{0}$, pretrained model $p_{0}(\boldsymbol{\theta})$, Unlabeled test data from different domain $\mathcal{U}_{1:T}$
[…]
3:for Domain shift $t=1$ to $T$ do
4:  for Test data $x\sim\mathcal{U}_{t}$ do
5:   Model predict for $x$ (Eq. (18))
6:   Update student model using $x$ (Eq. (16))
7:   Update teacher model via EMA (Eq. (17))
8:  end for
9:end for
@@ IV-D The algorithm
We illustrate the whole algorithm in Algorithm 2. We first transform an off-the-shelf pretrained model into BNN via the variational warm-up strategy (Section IV-C1). After that, we obtain a BNN, and for each domain shift, we forward and adapt each test data point in an MT architecture. For a data point $x$, we first predict the class label using the mixture of the source model and the teacher model (Section IV-C4). Then, we update the student model using VI, where we use cross entropy to compute the entropy term and use the mixture of priors for the KL term (Section IV-C2). Finally, we update the BNN teacher model via EMA (Section IV-C3). The process is feasible for any test data without labels.
Then, we show the whole CTTA algorithmic workflow utilizing variational approximation in VCoTTA:
Before testing: First, we adopt a variational warm-up strategy to inject stochastic dynamics into the model before adaptation. Given the source dataset $\mathcal{D}_{0}$, we can use a variational approximation of $p(\boldsymbol{\theta}|\mathcal{D}_{0})$ as Eq. (15), where we use the pretrained deterministic model $p_{0}(\boldsymbol{\theta})$ as the prior distribution.
During testing: We have the domain shift during the testing time, and we need to update the BNN in an MT structure. Then, at the beginning of the test time, we set the prior in task $t$ as $p_{t}(\boldsymbol{\theta})=\alpha\cdot p_{1}(\boldsymbol{\theta})+(1-\alpha)\cdot\bar{p}_{t}(\boldsymbol{\theta})$ for the variational approximation, where $p_{1}(\boldsymbol{\theta})\approx q_{0}(\boldsymbol{\theta})$ and $\bar{p}_{t}(\boldsymbol{\theta})\approx\bar{q}_{t}(\boldsymbol{\theta})$. $\bar{q}_{t}(\boldsymbol{\theta})$ means the real-time posterior of the teacher model for the $t$-th test domain, is constantly updated by $q_{t}(\boldsymbol{\theta})$ via EMA (see Eq. (17)). Note that we do not have $\bar{q}_{t}(\boldsymbol{\theta})$ for the first update in the $t$-th task, and use $q_{t-1}(\boldsymbol{\theta})$ construct the prior. This process is not required to inform the model that the domain produces a shift. Then, with the approximation of the prior $p_{t}(\boldsymbol{\theta})$ , we get $q_{t}(\boldsymbol{\theta})$ for student model at the test domain $t$ as Eq. (5). As a result, we recursively derive the following prior and posterior during the test time, thereby achieving the goal of VCoTTA.
IV-E Comparisons with Existing Methods
VCL [nguyen2018variational] is a classic continual learning (CL) study that uses VI. VCL inspires our work but has the following differences. First, VCL studies the supervised CL task, while our VCoTTA studies the unsupervised CTTA task. Second, CL only suffers from catastrophic forgetting (CF), while CTTA suffers from both CF and error accumulation. Third, to conduct BI, one needs to compute prior and likelihood. For the prior, the current prior of VCL is set to be the previous posterior, while in CTTA such a prior may be unreliable. For the likelihood, VCL can directly compute likelihood, CTTA is under the unsupervised setting, thus in our work, we deduce the BI in CTTA using conditional entropy. Last, to reduce error accumulation in the unsupervised scenario, we employ a mean-teacher update strategy using VI for the student model and EMA for the teacher model and compute a prior mixture to guide the student update. Moreover, VCL maintains an extra coreset from the training set, while VCoTTA never stores any data during the test time.
PETAL [brahma2023probabilistic] also estimates uncertainties in CTTA via BI. The BI formulation is similar between PETAL and ours, which is derived from [grandvalet2004semi], but PETAL uses a different method to conduct the inference. First, PETAL only does not estimate the model uncertainties via BNN. Second, PETAL ignores the unreliable prior in CTTA, and follows the VCL setting that still uses the previous posterior as the current prior. Third, we conduct BI using VI while PETAL use SWAG [maddox2019simple]. SWAG has advantages in terms of computational efficiency during training, but might not handle unreliable priors as effectively as VI since it doesn’t explicitly model the posterior distribution.
V Discussion and Theory Analysis
V-A BI during continual testing time
In this subsection, we derive the BI representation of CTTA in Eq. (4). First of all, the goal of CTTA is to learn a posterior distribution $p(\boldsymbol{\theta}|\mathcal{U}_{1:T}\cup\mathcal{D}_{0})$ from a source dataset $\mathcal{D}_{0}$, and a sequence of unlabeled test data from $\mathcal{U}_{1}$ to $\mathcal{U}_{T}$. Assuming the source dataset $\mathcal{D}_{0}$ is drawn from a distribution $\phi$, and $\tilde{\phi}_{t}$ is the $t$-th shift of the unlabeled test data. Let the parameters of the model be $\boldsymbol{\theta}$, then following the semi-supervised learning framework [grandvalet2004semi], we incorporate all input-generating distributions into the belief over the model parameters $\boldsymbol{\theta}$ as follows
$p(\boldsymbol{\theta}|\phi,\tilde{\phi}_{1},\cdots,\tilde{\phi}_{T})\propto$ $p(\boldsymbol{\theta})\exp\left(-\lambda_{0}H_{\boldsymbol{\theta},\phi}(Y|{X})\right)$ (19)
$\prod_{t=1}^{T}\exp\left(-\lambda_{t}H_{\boldsymbol{\theta},\tilde{\phi}_{t}}(Y|{X})\right),$
where the inputs $X$ are sampled i.i.d. from a generative model with parameters $\phi$, while the corresponding labels $Y$ are sampled from a conditional distribution $p(Y|X,\boldsymbol{\theta})$, which is parameterized by the model parameters $\boldsymbol{\theta}$. $p(\boldsymbol{\theta})$ is the prior distribution over $\boldsymbol{\theta}$, and $\{\lambda_{0},\lambda_{1},\cdots,\lambda_{T}\}$ are the factors for approximation weighting.
In Eq. (19), the term $H_{\boldsymbol{\theta},\phi}(Y|\boldsymbol{X})$ represents the cross entropy of the supervised learning, and the term $H_{\boldsymbol{\theta},\tilde{\phi}_{t}}(Y|\boldsymbol{X})$ for $t>0$ denotes the Shannon entropy of the unsupervised learning. Following [zhou2021training], we can empirically use a point estimation to get a plug-in Bayesian approach to approximate the above formula:
$p(\boldsymbol{\theta}|\mathcal{U}_{1:T}\cup\mathcal{D}_{0})$ (20)
$\propto$ $p(\boldsymbol{\theta})\prod_{\forall x,y\in\mathcal{D}_{0}}p(y|x,\boldsymbol{\theta})\exp\left(-\frac{\lambda_{0}}{|\mathcal{D}_{0}|}\sum_{\forall x\in\mathcal{D}_{0}}{H}(y|{x},\boldsymbol{\theta})\right)$
$\prod_{t=1}^{T}\exp\left(-\frac{\lambda_{t}}{|\mathcal{U}_{t}|}\sum_{\forall x\in\mathcal{U}_{t}}{H}(y|{x},\boldsymbol{\theta})\right).$
To make the formula feasible to CTTA, that is, no source data is available at the test time, we set $\lambda_{0}=0$. Since the source knowledge can be represented by $p(\boldsymbol{\theta}|\mathcal{D}_{0})\propto p(\boldsymbol{\theta})\prod_{\forall x,y\in\mathcal{D}_{0}}p(y|x,\boldsymbol{\theta})$, for the $t$-th test domain, the Bayesian inference in CTTA can be represented as follows:
$p(\boldsymbol{\theta}|\mathcal{U}_{1:t}\cup\mathcal{D}_{0})$ (21)
$\propto$ $p(\boldsymbol{\theta}|\mathcal{D}_{0})\prod_{i=1}^{t}\exp\left(-\frac{\lambda_{i}}{|\mathcal{U}_{i}|}\sum_{\forall x\in\mathcal{U}_{i}}{H}(y|{x},\boldsymbol{\theta})\right)$
$p(\boldsymbol{\theta}|\mathcal{U}_{1:t}\cup\mathcal{D}_{0})$ $\propto p(\boldsymbol{\theta}|\mathcal{U}_{1:t-1}\cup\mathcal{D}_{0})e^{-\lambda H(\mathcal{U}_{t}|\boldsymbol{\theta})}$ (22)
$=p_{t}(\boldsymbol{\theta})e^{-\lambda H(\mathcal{U}_{t}|\boldsymbol{\theta})},$
V-B Upper Bound of the Mixture of Two KL Divergencies
We refer to the lemma that was stated for the mixture of Gaussian in [singer1998batch].
Lemma 1
The KL divergence between mixture distributions $p=\sum_{i=1}^{k}\alpha_{i}p_{i}$ and $p^{\prime}=\sum_{i=1}^{k}\alpha_{i}p^{\prime}_{i}$ is upper-bounded by
Proof 1
By the log-sum inequality [cover1999elements], we have
$\leq$ $\int\sum_{i=1}^{k}\alpha_{i}p_{i}\log\cfrac{\alpha_{i}p_{i}}{\alpha_{i}p^{\prime}_{i}}=\sum_{i=1}^{k}\alpha_{i}\left(\int p_{i}\log\cfrac{\alpha_{i}}{\alpha_{i}^{\prime}}+\int p_{i}\log\cfrac{p_{i}}{p^{\prime}_{i}}\right)$
So the lower bound in Eq. (7) can be redefined as
$\text{ELBO}=$ $-\lambda\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta})}H(\mathcal{U}_{t}|\boldsymbol{\theta})-\mathrm{KL}\left(q(\boldsymbol{\theta})\parallel p_{t}(\boldsymbol{\theta})\right)$ (25)
$\geq$ $-\lambda\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta})}H(\mathcal{U}_{t}|\boldsymbol{\theta})-\alpha\cdot\text{KL}\left(q||{p}_{1}\right)$
$-(1-\alpha)\cdot\text{KL}\left(q||\bar{p}_{t}\right).$
Then, we have obtained a lower bound that can be optimized as the source prior distribution $q_{0}(\boldsymbol{\theta})$ and the teacher prior distribution $\bar{q}_{t}(\boldsymbol{\theta})$ are multivariate Gaussian distributions. Specifically, when the priors are multivariate Gaussian distributions, the KL term can be computed in closed form as
$\mathrm{KL}\left(\mathcal{N}(\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1})\parallel\mathcal{N}(\boldsymbol{\mu}_{2},\boldsymbol{\Sigma}_{2})\right)$ (26)
$=$ $\frac{1}{2}\Bigg(\text{tr}(\boldsymbol{\Sigma}_{2}^{-1}\boldsymbol{\Sigma}_{1})+(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})^{\top}\boldsymbol{\Sigma}_{2}^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})$
$-k+\ln\left(\frac{\det(\boldsymbol{\Sigma}_{2})}{\det(\boldsymbol{\Sigma}_{1})}\right)\Bigg),$
where $\boldsymbol{\Sigma}=\text{diag}(\boldsymbol{\sigma^{2}})$, $k$ represents the dimensionality of the distributions, $\text{tr}(\cdot)$ denotes the trace of a matrix, and $\det(\cdot)$ stands for the determinant of a matrix.
V-C Advantage of the Mixture of Gaussian Prior
In this subsection, we illustrate why the mixture of Gaussian prior are beneficial to CTTA. We first define what is a better distribution for CTTA. Assume there exists an ideal prior distribution $\hat{p}_{t}$, which effectively represents the distribution of the model after learning all past knowledge, including that from the source and unlabeled datasets. Then we can use the difference between a distribution and the ideal distribution $\hat{p}_{t}$ (here we use KL divergence) to measure the goodness of a distribution, i.e., $\text{KL}(\cdot||\hat{p}_{t})$.
Generally, neither the source prior $p_{1}$ (trained on labeled data) nor the adapted prior $\bar{p}_{t}$ (adapt on unlabeled data, being unreliable) can be completely consistent with $\hat{p}_{t}$. Considering that, as $t$ increases, the difference between $\bar{p}_{t}$ and $\hat{p}_{t}$ will increase without an upper bound due to the error accumulation (since $t$ is infinitely growing). The source prior $p_{1}$ cannot adapt to the unlabeled data, but it contains important information from the labeled data, and the ideal distribution cannot forget the source information too much, so we assume that the difference between $p_{1}$ and $\hat{p}_{t}$ is a constant, i.e., $\text{KL}(p_{1}||\hat{p}_{t})<U$, where $U$ is a constant upper bound. Accordingly, it can be considered that mixing the source prior $p_{1}$ and the adapted prior $\bar{p}_{t}$ in some way is beneficial for reducing $\text{KL}(\cdot||\hat{p}_{t})$.
In our method, we use Gaussian mixture $p_{t}=\alpha_{t}p_{1}+(1-\alpha_{t})\bar{p}_{t}$, where $\alpha$ is computed by Eq. (11). It is easy to illustrate the benefits of this idea using the following inequality:
$\text{KL}(p_{t}||\hat{p}_{t})$ $=\text{KL}\left[(\alpha_{t}p_{1}+(1-\alpha_{t})\bar{p}_{t})||\hat{p}_{t}\right]$ (27)
$\leq\alpha_{t}\text{KL}(p_{1}||\hat{p}_{t})+(1-\alpha_{t})\text{KL}(\bar{p}_{t}||\hat{p}_{t})$
$\leq\alpha_{t}U+(1-\alpha_{t})\text{KL}(\bar{p}_{t}||\hat{p}_{t}).$
If $\text{KL}(\bar{p}_{t}||\hat{p}_{t})\geq U$, which can be satisfied as mentioned above, then we have $\text{KL}(p_{t}||\hat{p}_{t})\leq\text{KL}(\bar{p}_{t}||\hat{p}_{t})$. This indicates that the mixed distribution $p_{t}$ is closer to the ideal distribution $\hat{p}_{t}$ than the adapted prior $\bar{p}_{t}$. A similar idea can be found in the stochastic restoration in CoTTA [wang2022continual], which randomly restores parts of the parameters of the current model into the parameters of the source model.
VI Experiment
VI-A Experimental Setting
VI-A1 Dataset